Math 524  University of South Carolina
D. Meade  Fall 1995

Homework Problems for Chapter 1 Unconstrained Optimization via Calculus

Section Page Problems
1.1  31 # 1abcd
1.2  31 # 3abc, 4b, 18
1.3  31 # 2ade, 6, 7abc
1.4  31 # 12abcdef, 13, 15
1.5  31 # 2ade, 8abc, 21, 33

\underline{Notes}

• Homework problems will be collected each Monday. In general, problems for each section are due on the first Monday following their discussion in class.

• Do not wait until the last minute to look at the problems. In fact, I suggest that you look at the problems in advance of our discussion of the material in class.

• I have tried to keep the number of problems to a minimum. However, some of these problems will require some thought. You should expect that some parts of them do not make sense the first time you read them. However, as you re-read the problem, and the text, you should find that the questions start to become clearer.

• I expect you to come to me for help whenever it’s needed. I believe it is very important (for this course) that you successfully work all these problems. My e-mail address is meade@math.sc.edu; my office phone is x6183.